### v1

# Collective emission and selective-radiance in atomic clouds and arrays coupled to a microring resonator

Abstract
We theoretically investigate the collective dipole-dipole interactions in atoms coupled to a nanophotonic microring resonator. The atoms can interact with each other through light-induced dipole-dipole interactions mediated by free space and through the resonator whispering-gallery modes. The differing characteristics and mismatched wavenumbers of these modes give rise to complex dynamics and provide new opportunities for controlling light–matter interactions. We explore these phenomena in the context of an experimentally realized atom cloud and study the potential of the proposed sub-wavelength atom arrays.
pacs:
I Introduction
The study of collective effects due to dipole-dipole interactions has been a major topic of interest since Dicke described the concept of superradiance in his seminal paper in 1954 Dicke (1954). There has been abundant research following this on the different aspects of collective dipole interactions like superradiance, subradiance, and collective Lamb shifts Bettles et al. (2015); Rehler and Eberly (1971); Friedberg et al. (1973); Gross et al. (1976); Scully (2009); Meir et al. (2014); Pellegrino et al. (2014); Browaeys et al. (2016); Bromley et al. (2016); Plankensteiner et al. (2017); Jennewein et al. (2018). These effects have been utilized for applications in quantum control, photon storage, and quantum information Bettles et al. (2016); Facchinetti et al. (2016); Shahmoon et al. (2017); Manzoni et al. (2018).
Great progress has been achieved in engineering nanophotonic interfaces for versatile control over light-matter interaction. The interaction between atoms has been experimentally implemented using nanofibers Solano et al. (2017); Le Kien and Rauschenbeutel (2017); Pennetta et al. (2022a, b); Liedl et al. (2023, 2024); Bach et al. (2024); Su et al. (2023), photonic crystals Goban et al. (2015); Hood et al. (2016), and slot waveguides Ritter et al. (2018); Alaeian et al. (2024). Since the atoms are usually trapped close to dielectric surfaces, the complexities of modeling the interactions have been explored in Refs. Asenjo-Garcia et al. (2017); Jones et al. (2018); Svendsen and Olmos (2023); Cardenas-Lopez et al. (2023).
Recently, neutral atoms have been laser-cooled and trapped adjacent to a nanophotonic microring resonator Zhou et al. (2024) to be used as a versatile light-matter interaction platform. The ring-resonator can act as a whispering gallery mode (WGM) cavity and facilitate chiral atom-light interactions.
This platform has several unique properties. The atoms can interact with each other through two different types of interaction, through free space and through the resonator, each with a different characteristic wavenumber. The interaction strength with the resonator can also be varied by adjusting the position of the atoms.
One of the potentials of such atom-nanophotonic interface is the concept of "selective radiance" as was described in Ref. Asenjo-Garcia et al. (2017). These selectively radiant states can be simultaneously superradiant to a desired guided mode and subradiant to undesired or error-inducing modes. Due to the nature of the coupling with the ring resonator, the atoms will superradiantly couple with the resonator, and the decay rate will scale with the atom number Zhou et al. (2024). If the decay into free space can be reduced due to the collective dipole-dipole interactions, the interface becomes more robust against error from spontaneous emission into the vacuum.
This light-matter platform could have potential applications involving quantum memories Gorshkov et al. (2007), quantum simulation Gonzalez-Tudela et al. (2011); Ramos et al. (2014); Pichler et al. (2015); Ramos et al. (2016), quantum networks Stannigel et al. (2012), and applications in quantum computing Paulisch et al. (2016); González-Tudela et al. (2015, 2017). It also holds the potential for investigating photon interactions Roy et al. (2017); Zhang and Mølmer (2019); Sheremet et al. (2023), non-Markovian effects Zhang (2023), and topological models Zhang and Mølmer (2020).
In this paper, we theoretically and computationally model the atom and ring resonator system described in Ref. Zhou et al. (2025) to understand the physics underlying the interactions and explore further possibilities. We give particular emphasis on how the collective dipole-dipole interactions mediated by free space modes can affect the interaction with the microring resonator. We describe the system and the numerical methods used in Sec. II. We simulate the current experiment, which involves a trapped atomic cloud above the ring resonator in Sec. III.2. In Sec. III.3, we explore the potential of an atom array near the ring resonator and consider the effect of disorder that might arise in experiments. This work serves as a companion paper to Ref. Zhou et al. (2025), which describes the experimental aspects and discusses the measured photon decay rates.
II Methods
The experimental system consists of a microring resonator coupled to a bus waveguide, as represented in Fig. 1. The external coupling rate of the resonator to the bus waveguide ($\kappa_{e}$) is comparable to the intrinsic loss rate of the resonator ($\kappa_{i}$), which makes the resonator close to the critical coupling condition Zhou et al. (2023). The microring resonator acts as a whispering gallery mode cavity that can sustain light circulating in the two directions.
Figure 1: Schematic of the experimental system. (a) Cold atoms are trapped above the microring resonator, interacting with a WGM. The WGM is circularly polarized to drive the $\sigma^{+}$ cycling transition |$\ket{g}\rightarrow\ket{e}$. (b,c) Interacting atomic ensemble and organized atom array collectively emit photons into the WGM and the free space modes with photon emission rates $R_{c}$ and $R_{f}$, respectively.
The experiment uses the fundamental Transverse Magnetic-like (TM) WGM. The electric field of the WGM is circularly polarized above the waveguide, and the polarization is locked to the direction of circulation in the resonator. The atoms are polarized to a stretched state and couples to the resonator mode via a cycling transition. As such, the atoms are constrained to emit into the same mode with which it is excited. This configuration results in the atom-light interaction through the waveguide being chiral. We only consider interaction with a fixed direction of circulation, as in Fig. 1(a).
Cold atoms are cooled and trapped above the microring resonator with a peak density of around $10^{13}\,\mathrm{cm^{-3}}$ in an almost cigar-shaped configuration. The number of atoms in the cloud can be varied from a few to up to 60 atoms. These atoms are coupled to the evanescent field of the circulating WGM in the resonator. Since the evanescent field decreases exponentially above the ring, each atom has a different interaction strength with the WGM. This interaction strength can also be controlled by moving the atom cloud closer or farther away from the resonator.
To model this system, we use the master equation to evolve the density matrix of the system given by
$\frac{d\hat{\rho}}{dt}=-\frac{i}{\hbar}[\hat{H}_{eff},\hat{\rho}]+\mathcal{L}(\hat{\rho}),$ (1)
where $\hbar$ is the reduced Planck constant, $\hat{\rho}$ is the density matrix of the system, $\hat{H}_{eff}$ is the effective Hamiltonian given by Eq. (15) and $\mathcal{L}(\hat{\rho})$ is the Lindblad super-operator which describes the lossy interaction, as defined in Eq. (16).
Since we are restricted to the low-intensity limit, there can be only one excitation in the system at any time. We consider $N$ atoms in the ensemble and the Hilbert space of the atom excitations can be simplified to $N+1$ states. The raising and lowering operators of the $j^{\mathrm{th}}$ atom are represented by $\hat{\sigma}^{+}_{j}$ and $\hat{\sigma}^{-}_{j}$ respectively. The state in which all atoms are in the ground state is represented by $|g\rangle$, and the states where only the $j^{\mathrm{th}}$ atom is excited will be represented by $|e_{j}\rangle=\hat{\sigma}^{+}_{j}|g\rangle$. The single-atom decay rate of the atoms will be denoted by $\Gamma_{0}$ and will serve to set the timescale of the system.
Since there are two types of interactions, we first describe the collective dipole-dipole interactions mediated by the non-guided modes in free space, followed by the effect of coupling the atomic ensemble to the cavity.
II.1 Collective dipole-dipole interaction
Since the separations between adjacent atoms are small and comparable to the wavelength of light, there will be collective dipole-dipole interactions in the absence of the cavity. This can be modeled using the free space dyadic Green’s function $\mathbf{G}(\mathbf{r}_{i},\mathbf{r}_{j},\omega_{0})$, where $\omega_{0}$ is the resonant light frequency and $\mathbf{r}_{i}$, $\mathbf{r}_{j}$ are the position vectors of atom $i$ and $j$.
The dispersive and dissipative interaction strengths $J_{ij}$ and $\Gamma_{ij}$ are given by the Green’s function as
$J_{ij}$ $=-\frac{\mu_{0}\omega_{0}^{2}}{\hbar}\vec{d}^{*}\cdot\mathrm{Re}\{\mathbf{G}(\mathbf{r}_{i},\mathbf{r}_{j},\omega_{0})\}\cdot\vec{d}$ (2)
$\Gamma_{ij}$ $=\frac{2\mu_{0}\omega_{0}^{2}}{\hbar}\vec{d}^{*}\cdot\mathrm{Im}\{\mathbf{G}(\mathbf{r}_{i},\mathbf{r}_{j},\omega_{0})\}\cdot\vec{d},$ (3)
where $\mu_{0}$ is the vacuum permeability and $\vec{d}$ is the dipole operator of the atoms. The dispersive part of the Green’s function induces a resonant dipole-dipole exchange Hamiltonian,
$\hat{H}_{dd}^{f}=\hbar\sum_{i\neq j}J_{ij}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}$ (4)
and the dissipative part of the free space interaction is described by the Lindblad super-operator,
$\mathcal{L}_{f}(\hat{\rho})=\sum_{i,j}\frac{\Gamma_{ij}}{2}\big{[}2\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{j}^{+}-\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\big{]}.$ (5)
The interaction matrix associated with the free space Green’s function $G_{ij}^{f}$ can be written as
$G^{f}_{ij}=J_{ij}-i\frac{\Gamma_{ij}}{2}.$ (6)
The effects of the atoms being in the vicinity of the dielectrics are negligible in this case due to the distance ($z\gtrsim\lambda_{0}/2$, where $\lambda_{0}$ is the resonant wavelength) from the surface (Supplementary of Ref. Zhou et al. (2025)). Therefore, it would suffice to take the simple vacuum Green’s function to make effective qualitative calculations. The interaction matrix $G_{ij}^{f}$ will become
$\begin{split}G^{f}_{ij}=-i\frac{\Gamma_{0}}{2}\bigg{[}h_{0}^{(1)}&(kr_{ij})+\frac{3(\hat{r_{ij}}\cdot\hat{d})(\hat{r_{ij}}\cdot\hat{d}^{*})-1}{2}h_{2}^{(1)}(kr_{ij})\bigg{]},\end{split}$ (7)
where $\hat{r}_{ij}=\mathbf{r}_{ij}/r_{ij}$ is the unit vector along $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$, $r_{ij}=|\mathbf{r}_{ij}|$ is the norm, and $h_{l}^{(1)}$ are the outgoing spherical Hankel function of angular momentum $l$. $h_{0}^{(1)}(x)=e^{ix}/(ix)$ and $h_{2}^{(1)}(x)=e^{ix}(-3i/x^{3}-3/x^{2}+i/x)$. The coefficient in front of $h_{2}^{(1)}$ depends on the orientation of the dipoles and the driven transitions, which will be $(1-3\cos^{2}{\theta})/4$ for $\sigma^{+}$ circularly polarized light quantized in the x-direction.
II.2 Interaction with the cavity
Once the atomic ensemble is brought to the vicinity of the microring resonator, the atoms can couple with the evanescent field of the cavity and interact. We consider a single mode cavity supporting the $\sigma^{+}$ circularly polarized mode. This cavity interaction can be modeled as a harmonic oscillator with raising (lowering) operators, $\hat{a}^{\dagger}(\hat{a})$. The Hamiltonian of the atom-cavity interaction can be represented as
$\begin{split}\hat{H}_{I}=&\sum_{j}g_{j}(e^{-i\phi_{j}}\hat{a}\hat{\sigma}^{+}_{j}+e^{i\phi_{j}}\hat{a}^{\dagger}\hat{\sigma}^{-}_{j})\\ &-\sum_{j}\Delta_{A}\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}+\eta(\hat{a}+\hat{a}^{\dagger})-\Delta_{C}\hat{a}^{\dagger}\hat{a},\end{split}$ (8)
where $g_{j}$ depends on the evanescent electric field at the position of atom $j$, $\Delta_{A}=\omega-\omega_{A}$ and $\Delta_{C}=\omega-\omega_{C}$ are the detuning of the atomic resonance and the cavity resonance from the input light frequency, respectively. $\eta$ is the classical driving rate due to the external driving through the bus waveguide. $\phi_{j}=\mathbf{k}\cdot\mathbf{r}_{j}$, where $\mathbf{k}$ is the wavevector of the ring resonator mode at position $\mathbf{r}_{j}$. Here, $|\mathbf{k}|=k_{\mathrm{wg}}$ is the wavenumber of the waveguide mode; $k_{\mathrm{wg}}>k_{0}$, the wavenumber associated with the resonant wavelength of light.
In the experiment, single-atom cooperativity plays a more relevant role in describing the strength of the atom-cavity interaction. The single atom cooperativity of atoms is given by,
$C=\frac{4g_{j}^{2}}{(\kappa_{i}+\kappa_{e})\Gamma_{0}},$ (9)
which depends on $g_{j}$ and in turn depends on the position of each atom. It decays exponentially as the distance between the atom and the resonator surface increases. Given an atomic density distribution, the average single-atom cooperativity will be denoted as $C_{1}$. In general, the total cooperativity of the waveguide interaction scales as $NC_{1}$.
The microring resonator is in the bad-cavity limit, where the cavity dynamics happen at a much faster rate than the dynamics of the atomic internal states ($\kappa_{e},\kappa_{i}\approx 100\Gamma_{0}$). The cavity field can be adiabatically eliminated to more intuitively capture the dynamics of the internal excitations of the atoms. This will also drastically decrease the computational requirements. By enforcing that the cavity reaches a steady state at every time-step of the internal dynamic evolution, we can trace out the cavity modes and obtain a simplified model to study the interactions between the atoms. Further details on the adiabatic eliminations is explained in Appendix A.
After eliminating the cavity, the interaction of the atoms through the cavity will have a form similar to the collective dipole interactions described in Sec. II.1. There will be a dipole-dipole exchange Hamiltonian and a loss-inducing Lindblad operator
$\hat{H}_{dd}^{c}=\hbar\sum_{i\neq j}\mathrm{Re}\{G_{ij}^{c}\}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}$ (10)
$\mathcal{L}_{c}(\hat{\rho})=\sum_{i,j}\big{[}2iG_{ij}^{c}\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{j}^{+}+\mathrm{Im}\{G_{ij}^{c}\}(\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}+\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-})\big{]},$ (11)
where the interaction matrix through the cavity $G_{ij}^{c}$ will be given by
$\begin{split}G^{c}_{ij}=\frac{g_{i}g_{j}^{*}}{\tilde{\kappa}}e^{-i(\phi_{i}-\phi_{j})},\end{split}$ (12)
and $\tilde{\kappa}=\Delta_{C}+i(\kappa_{i}+\kappa_{e})/2$.
The driving of the system through the bus waveguide will appear in a form similar to a rotating wave laser Hamiltonian
$\hat{H}_{L}=\hbar\sum_{j}\left[-\Delta_{A}\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}+\frac{\Omega_{j}}{2}\hat{\sigma}_{j}^{+}+\frac{\Omega^{*}_{j}}{2}\hat{\sigma}_{j}^{-}\right],$ (13)
where $\Omega_{j}$ is the coupling with which each atom is excited due to the evanescent mode of the microring. It can be connected with the classical driving rate $\eta$ as
$\Omega_{j}=\frac{g_{j}e^{i\phi_{j}}}{\tilde{\kappa}}\eta.$ (14)
To describe the complete evolution of the system after adiabatic elimination, the effective Hamiltonian and total Lindblad operator in Eq. (1) will be given by
$\hat{H}_{eff}=\hat{H}_{dd}^{f}+\hat{H}_{dd}^{c}+\hat{H}_{L}$ (15)
$\mathcal{L}(\hat{\rho})=\mathcal{L}_{f}(\hat{\rho})+\mathcal{L}_{c}(\hat{\rho}).$ (16)
The density matrix can then be time-evolved using the master equation shown in Eq. (1) with Runge-Kutta methods. The full cavity model and the adiabatically eliminated Lindblad model have both been numerically compared and verified in the single excitation limit. The limitation to adiabatic elimination is that the cavity cannot be in the strong-coupling regime.
An important approximation is that the atoms are stationary within the time scale of the internal excitation decay. This is reasonable since the atoms in the cloud have a mean velocity of 3 cm/s, which corresponds to a couple of nanometers of movement in one excitation lifetime. Hence, to capture the dynamics of the ensemble, we can use a Monte Carlo method to take the average of many different possible positions for the atoms. In each iteration, the positions of the atoms are randomly assigned according to the configuration of the atom cloud.
II.3 Mean field equations
Time evolving the master equation in this particular scenario can lead to a few complications in the calculations. Since the atom cloud is considerably dense, the atoms in the random positions could end up being relatively close to another atom. This will cause the free space Green’s function to increase significantly, beyond the energy scale of the calculation and requiring prohibitively small time-steps. Additionally, since these systems tend to have some states with subradiant character, the time to reach the steady state is much longer than the timescale of typical excitation decay. These issues make it challenging to reach convergence when using more than a few atoms. Hence, we utilize the weak field limit and can derive a procedure to directly calculate the final state expectation values using eigenmode decomposition.
The equation of motion of the expectation values at low intensities can be written as
$\begin{split}\frac{d\langle\hat{\sigma}^{-}_{i}\rangle}{dt}=\quad i\Delta_{a}\langle\hat{\sigma}^{-}_{i}\rangle+i\Omega_{i}-i\sum_{j}(G_{ij}^{c}+G_{ij}^{f})\langle\hat{\sigma}^{-}_{j}\rangle,\end{split}$ (17)
which can be simplified as
$\dot{\vec{\sigma}}=i\mathbf{M}\vec{\sigma}+i\vec{\Omega},$ (18)
where $\vec{\sigma}=\{\langle\hat{\sigma}^{-}_{1}\rangle,...,\langle\hat{\sigma}^{-}_{N}\rangle\}$, $\vec{\Omega}=\{\Omega_{1},...,\Omega_{N}\}$, and $\mathbf{M}=\Delta_{A}\mathbb{1}-\mathbf{G}_{c}-\mathbf{G}_{f}$ is the combined coupling matrix of the two types of interactions.
II.4 Eigenmodes of the excitation
To understand the characteristics of the system, we can study the eigenmodes of the Green’s function of the two different interactions. Apart from having a mismatched wavenumber, the nature of the interaction matrices are different.
For resonant probe and atoms resonant with cavity, $\Delta_{\mathrm{a}}=\Delta_{\mathrm{c}}=0$, $\mathbf{G}_{c}$ is a skew-Hermitian matrix ($\mathbf{G}_{c}^{\dagger}=-\mathbf{G}_{c}$), which results in purely imaginary eigenvalues. There is a single non-zero eigenmode with decay rate $NC_{1}\Gamma_{0}$ with the rest $N-1$ being completely dark subradiant states. In this case, the contribution of $\mathbf{G}_{c}$ to the evolution of atomic coherence $\vec{\sigma}$ will be purely dissipative, which is expected for a resonant atom-cavity interaction. $\mathbf{G}_{f}$, on the other hand, is a complex symmetric matrix ($\mathbf{G}_{f}^{\mathrm{T}}=\mathbf{G}_{f}$) and is both dissipative and dispersive. The combined coupling matrix $\mathbf{M}$ can be diagonalized with N eigenvectors $\vec{v}_{\alpha}$ of complex eigenvalues
$\lambda_{\alpha}=J_{\alpha}-i\frac{\Gamma_{\alpha}}{2}.$ (19)
Here, the $\vec{v}_{\alpha}$ will be the right eigenvectors of the complex matrix $\mathbf{M}$, and further details about the orthogonality relations are explained in Appendix B.
An atomic state can be decomposed using the eigenvectors as
$\vec{\sigma}=\sum_{\alpha}w_{\alpha}\vec{v}_{\alpha},$ (20)
where $w_{\alpha}$ is the weight associated with eigenmode $\alpha$.
When driven to the steady state (SS) using a long excitation pulse, the final state $\vec{\sigma}_{\mathrm{SS}}=\sum_{\alpha}w_{\alpha}^{\mathrm{SS}}\vec{v}_{\alpha}$ can be calculated using $\dot{\vec{\sigma}}=0$. In the case of a short pulse, the state can be directly initialized into a timed-Dicke state (TDS) with a small excitation population since the pulse time will be too short to cause mixing into any other states, $\vec{\sigma}_{\mathrm{TDS}}=\sum_{\alpha}w_{\alpha}^{\mathrm{TDS}}\vec{v}_{\alpha}\propto\vec{\Omega}$. The method to calculate $w_{\alpha}^{\mathrm{SS}}$ and $w_{\alpha}^{\mathrm{TDS}}$ is described in Appendix B.
Using these methods to directly calculate the final state of the system instead of time propagating the density matrix greatly decreases the computation time. The convergence of the results from the mean field, adiabatic elimination, and the full density matrix has been numerically verified in the weak field limit.
II.5 Photon emission into the cavity and free space
We study the evolution of the system right after the excitation pulse is switched off. The system will then evolve under the equation $\dot{\vec{\sigma}}=i\mathbf{M}\vec{\sigma}$. For an initial state $\vec{\sigma}_{0}=\sum_{\alpha}w_{\alpha}\vec{v}_{\alpha}$ at $t=0$, the evolution will be $\vec{\sigma}(t)=\sum_{\alpha}w_{\alpha}e^{i\lambda_{\alpha}t}\vec{v}_{\alpha}$. Atomic excitation will decay to the ground state and emit photons into different modes. The rate of photon emission can be calculated from the total de-excitation rate in the system.,
$\begin{split}R&=-\langle\dot{\hat{e}}\rangle=-\frac{d(\vec{\sigma}^{\dagger}\vec{\sigma})}{dt}=\vec{\sigma}^{\dagger}(i\mathbf{M}^{\dagger}-i\mathbf{M})\vec{\sigma}\\ &=2i\vec{\sigma}^{\dagger}\mathbf{G}_{c}\vec{\sigma}-2i\vec{\sigma}^{\dagger}\mathrm{Im}\{\mathbf{G}_{f}\}\vec{\sigma},\end{split}$ (21)
where $\langle\hat{e}\rangle$ is the total excitation probability in the system and we have used $\mathbf{G}_{c}^{\dagger}=-\mathbf{G}_{c}$ and $\mathbf{G}_{f}^{T}=\mathbf{G}_{f}$. The rate of photons emitted into the cavity mode and the free space modes can be separately calculated as
$R_{c}=2i\vec{\sigma}^{\dagger}\mathbf{G}_{c}\vec{\sigma}$ (22)
$R_{f}=-2\vec{\sigma}^{\dagger}\mathrm{Im}\{\mathbf{G}_{f}\}\vec{\sigma}.$ (23)
Experimentally measured transmission rate at the output of the bus waveguide is directly proportional to the photon emission rate $R_{c}$. Photon emission into free space is typically collected by optics spanning over a finite solid angle. If the angular pattern of photon emission is fixed in time, the measured photon rate is proportional to $R_{f}$.
Although the amplitude associated with each eigenstate decays exponentially in time, the emission rates contain the cross-coupling between all the eigenmodes, leading to more complicated time evolution dynamics. Hence, the rates, as well as the total excitation probability, do not decay purely exponentially. The decay rate of the light emitted varies over time. In this paper, the focus will be on the instantaneous decay rate immediately after the drive is turned off in order to study the initial time behaviors of the system. The decay rates of the excitation into the free space and cavity modes have been defined as
$\gamma_{f}=\frac{R_{f}}{\langle\hat{e}\rangle}\quad\mathrm{and}\quad\gamma_{c}=\frac{R_{c}}{\langle\hat{e}\rangle}.$ (24)
The former is approximately independent of the single atom cooperativity $C_{1}$. This is because, at low intensities, the decay rate into free space is solely dependent on the relative phase and the relative amplitude of the excitation between the atoms. While shifting the mean atomic position in the z-axis and changing the atom-cavity interaction strength could change the absolute excitation probability, it does not change the relative amplitude and phase between the atoms.
II.6 Decay rates of photon emission
The tools available in the experimental setup limits the reliable measurement of the photon decay rate to the cavity ($\dot{R}_{c}/R_{c}$). The photon emission rate into free space was not measured Zhou et al. (2024). In systems with exponential decay, this measurement is equivalent to the decay rate of the excitation in the system. However, this measurement and the decay rates defined in Eq. (24) are not equivalent for non-exponential decay dynamics. Reliably measuring the excitation decay rate, like $\gamma_{c}$ or $\gamma_{f}$, remains a challenge in most experiments. In this particular platform, where the interaction strength can be varied and non-exponential decay is present, the discrepancy can be larger. Hence, the experimental paper in Ref. Zhou et al. (2025) operationally defines the decay of the photon emission rate into the free space and cavity modes as
$\Gamma_{f}=\frac{-\dot{R}_{f}}{R}\quad\mathrm{and}\quad\Gamma_{c}=\frac{-\dot{R}_{c}}{R},$ (25)
where $R=R_{f}+R_{c}$ is the total photon emission rate. The measurement of the decay rate of $R_{c}$ can be related to these decay rates as
$\Gamma_{\mathrm{exp}}=-\frac{\dot{R}_{c}}{R_{c}}=\Gamma_{c}+\Gamma_{f}\theta,$ (26)
where $\theta=(\dot{R}_{c}/R_{c})/(\dot{R}_{f}/R_{f})$.
It is important to note that the difference between $\gamma_{c},\gamma_{f}$ and $\Gamma_{c},\Gamma_{f}$ vanishes in the case of single exponential decay. This is generally true for photon emission from an ordered atom array, where the excitation is largely confined to a single eigenmode.
III Results
We consider two main configurations: atoms in a randomly distributed cloud and an ordered array of atoms. The focus of this study will be on the decay characteristics due to emission into free space and into the cavity.
In the experiment, the most relevant observable is the time evolution of the photon emission rate into the cavity mode, which can be measured as a change in transmission rate from the output waveguide. These measurements can be utilized to verify the theoretical calculations. Further calculations can be performed to expand the scope and understanding of the experiment.
The atoms can be excited using either a short Gaussian pulse or a long-duration pulse that drives the system to the steady state. The collective excited states associated with these situations are considerably different. In the case of the short pulse, the atoms will get excited into a state similar to a spin-wave excitation known as the timed-Dicke state (TDS) Scully and Svidzinsky (2009) with a wavenumber corresponding to that of the resonator mode. Since the pulse time is short, the detuning will also play a negligible role. This will be a superradiant state of the atom-cavity interaction. For decay into free space, the TDS has a higher wavenumber than $k_{0}$, and is typically associated with subradiance.
The system can also be driven into the steady state (SS) by using a significantly longer pulse duration. Since there is constant driving, the system is usually pumped into excited states exhibiting slower decaying modes. The detuning also plays an important role in determining which modes are excited in this case.
III.1 Evanescent coupling
The atoms couple with the resonator mode through an evanescent coupling. The coupling decreases exponentially with distance from the resonator. The atoms’ average position along the z-axis depends on the trapping potential, allowing control of the single atom cooperativity $C_{1}$. This provides an opportunity to study the behavior of collective photon emissions as a function of the atom-cavity interaction strength.
The experiment controls the average position of the atoms to be $\geq 400\mathrm{nm}$ away from the surface, which corresponds to $C_{1}\lesssim 0.05$. In the proposed plan to implement an atom array, the trapping methodology can afford the atoms to be much closer to the resonator waveguide, leading to much higher average cooperativities ($C_{1}\sim 10$).
Although the position-dependent coupling can provide an opportunity for controlling the atom-WGM photon interaction strength, it can also cause issues when it comes to the randomized position of atoms in the cloud. The spread of the atoms in the z-direction due to the trapping potential will cause a stochastic non-uniformity in the single-atom cooperativity of the atoms. This causes a considerable amount of broadening and decoherence in the system.
III.2 Atom Cloud
Figure 2: (a,b) Ensemble averaged decay rate into free space when an atom cloud is excited to (a) the steady state (SS) or (b) the timed-Dicke state (TDS) with a WGM of wavenumber $k_{\mathrm{wg}}$. The color scale depicts the number of atoms in the atom cloud. (c,d) Histograms of the distribution of the excitation decay rate into (c) free space and (d) into the cavity when taking individual random realizations of the atomic cloud with $N=60$ atoms and $C_{1}=0.05$. The average value of the histogram has been plotted as vertical dashed lines of the corresponding color. The results are averages over 5000 random configurations.
Figure 3: The measured and calculated linewidth of the steady state transmission spectrum as a function of the number of atoms for $C_{1}=0.05$. Blue squares with error bars denote experimental data. Red circles with lines show theoretical calculations. The black triangles and orange diamonds with dashed lines show the theoretical calculation of hypothetical situations where there is no free space collective dipole-dipole interaction or when there are no stochastic effects
Figure 4: The ratio of decay rates for a cloud of atoms with uniform $C$ for the TDS and the SS. In both cases, the excitation decay ratio $\gamma_{c}/\gamma_{f}$ tends to be close to $NC_{1}$. The photon rate decay ratio $\Gamma_{c}/\Gamma_{f}$ increases or decreases depending on the type of excitation. The results are averaged over 2000 random configurations.
In the current state of the experiment, a cloud of atoms is trapped and cooled near the microring resonator in a trap with a cigar-shape of around $2\mu\mathrm{m}$ along the waveguide. In the transverse direction, the cloud is tightly trapped to an r.m.s. of around $100\mathrm{nm}$. The position dependence on the z-direction above the resonator is not Gaussian Zhou et al. (2024). It has an approximate r.m.s. of $430\mathrm{nm}$.
The system can be directly driven with a large wavenumber. Unlike in free space, where systems are conventionally only driven by the free space modes with wavenumber $k_{0}$, the effective refractive index of the resonator waveguide determines the wavenumber of the drive. Figure 2 (a) and (b) explore how driving an atomic cloud with a different wavenumber can affect the free space decay rate. The system is being excited using either short or long pulses, which can exhibit contrasting subradiant (SS) or superradiant (TDS) emission dynamics.
When driven using a short pulse, the system is excited into the TDS with the wavenumber $k\sim 1.7k_{0}$ While one might expect this state to be subradiant to free space, Ref. Grava et al. (2022) has shown that for an atom cloud, the smallest decay rate for such high-k states is $\Gamma_{0}$. This can be seen in Fig. 2 (b), where the free space decay rate $\gamma_{f}$ for all atom numbers converge to $\Gamma_{0}$ as the wavenumber increases beyond $k_{0}$. This happens because the dephasing from atoms being very close to each other in a random cloud acts against the longer-range coherence that is built up to cause subradiance. The TDS will be perfectly superradiant to the cavity with an average decay rate of $NC_{1}\Gamma_{0}$ while having a decay rate of $\Gamma_{0}$ into free space. Therefore, the TDS will have a total decay rate of $(1+NC_{1})\Gamma_{0}$.
When the atom cloud reaches the SS, the system displays slower photon emission dynamics. As shown in Fig. 2 (a), the free-space decay rate exhibits subradiance and the rate decreases with an increasing number of atoms. While subradiance can occur even when the system is excited with a free-space wavenumber $k_{0}$, the subradiant effect is enhanced when driven with higher wavenumbers $k>k_{0}$.
These effects can also be seen in Fig. 2 (c) and (d), which show the distribution of the decay rates into the free space and cavity modes without ensemble averaging. The free space decay rate ($\gamma_{f}$) is suppressed in the case of the SS while clustered around $\Gamma_{0}$ for the TDS. For the decay rate into the cavity ($\gamma_{c}$), the TDS has a large spread of decay rates, but the average is close to $NC_{1}\Gamma_{0}$, where $N$ is the atom number in the system. On the other hand, for the SS, $\gamma_{c}$ averages to around $20C_{1}\Gamma_{0}$, which is still superradiant but not as good as the with the TDS.
The large spread in $\gamma_{c}$ is primarily due to the stochastic non-uniformity of the single atom cooperativity $C$, as mentioned in Sec. III.1. When $C$ is uniform for all the atoms, the distribution of $\gamma_{c}$ will be a delta function at $NC_{1}\Gamma_{0}$ for the TDS and will have a considerably smaller spread for the SS. This emphasizes the importance of minimizing the spread in the z-position for systems that have evanescent coupling.
On the other hand, the steady state transmission spectrum through the output bus waveguide can be calculated and compared with experimental measurements. The linewidth can provide some indication of the the decay properties of the system. Figure 3 compares the experimentally measured and calculated linewidths of the spectrum, which are both determined using simple Lorentzian fits. The mean atom number $N$ in the experiment is extracted by fitting the steady-state transmission spectrum assuming a Gamma distribution of $NC$ with mean value $C_{N}$ and $N=C_{N}/C_{1}$ (see Ref. Zhou et al. (2024)).
Without the free space interaction, the system will be excited to only the superradiant mode of the atom-cavity interaction, and the linewidth will scale as $NC_{1}$. This is depicted by the black dashed line. In the presence of free space collective interaction, it facilitates the mixing between the superradiant and the completely dark subradiant states of the atom-cavity interaction. This pulls down the total linewidth as $N$ increases. This can be seen more clearly from the theoretical data points.
However, in practical implementations, the expected linewidth suppression is significantly weakened due to multiple stochastic effects. The primary contributor is the non-uniformity in single-atom cooperativity $C$ (see Sec. III.1), which broadens the linewidth. Additionally, the AC Stark shift from trapping fields varies with atomic position, introducing a slight spectral asymmetry and further broadening the linewidth. Moreover, the number of trapped atoms follows a Poissonian distribution, which introduces further broadening, especially for small $N$. The orange diamonds in Fig. 3 depict the hypothetical where these stochastic effects are absent. This linewidth shows a greater degree of suppression for larger $N$ and approaches $\Gamma_{0}$ when $N\rightarrow 1$.
Another aspect of the spectrum is the average line-shift. The atom-cavity Green’s function is purely dissipative and will not contribute to line-shifts. On the other hand, the interaction through free space will contribute to a shift depending on the density of the atoms. Therefore, a shift in the spectrum due to the free-space interaction can be expected. Contrary to expectations, the shifts observed experimentally were an order of magnitude higher than the calculations. Several effects, like the AC Stark shift, vicinity to a dielectric surface, and motion of the atoms were explored but were unable to explain the discrepancy. Much larger shifts are expected in the case of atom arrays, and clearer answers are expected once they are implemented in the next stage of the experiment.
The ratio of decay rates into the desired versus the undesired modes describes the effectiveness in mitigating photon emission into the undesired channels. Figure 4 shows the ratio of decay rates for the excitation ($\gamma_{c}/\gamma_{f}$) and the photon decay rate ($\Gamma_{c}/\Gamma_{f}$). In this plot, we assume a uniform cooperativity $C$ to simplify numerical convergence. Incorporating non-uniformity in $C$ does not lead to qualitative differences.
In the case of the TDS, $\gamma_{c}/\gamma_{f}\sim NC_{1}\Gamma_{0}$, showing the same scaling when the atoms are far apart and collective dipole interactions through free space are absent. This behavior is already seen in Fig. 2(c,d), where the average $\gamma_{c}/\Gamma_{0}\approx NC_{1}$ and $\gamma_{f}\approx\Gamma_{0}$. Surprisingly, in the case of the SS, although $\gamma_{f}$ is suppressed to be subradiant, $\gamma_{c}$ is also reduced to still give a scaling of $NC_{1}$ for the excitation decay ratio.
The photon rate decay ratio $\Gamma_{c}/\Gamma_{f}$ can exceed or fall below the expected $NC_{1}$ scaling depending on the excitation conditions. This happens because, although $\Gamma_{c}$ scales as $NC_{1}$ in both cases, $\Gamma_{f}$ increases or decreases with $N$ depending on the type of excitation. This can be seen in Fig. 4(c) in Zhou et al. (2025) where $\Gamma_{f}$ increases beyond $\Gamma_{0}$ with $N$ for the TDS and decreases below $\Gamma_{0}$ for the SS.
In the associated experimental paper Zhou et al. (2025), the ratio $\theta$ in Eq. (26) serves as an important factor to correlate experimentally measured decay rates to the actual decay rates of the system. To get an intuition behind this $\theta$, we can express it as $\theta=(\Gamma_{c}/\Gamma_{f})/(\gamma_{c}/\gamma_{f})$. In the case of the TDS, $\Gamma_{c}/\Gamma_{f}<NC_{1}$ resulting in the $\theta<1$, while for the SS, $\Gamma_{c}/\Gamma_{f}>NC_{1}$ resulting in the $\theta>1$. Further analysis on the implications of $\theta$ is explored in the Supplementary of Ref. Zhou et al. (2025).
III.3 Atom Arrays
Figure 5: The decay rate into free space when a perfect atom array with 20 atoms is excited to the TDS through the resonator, as the separation $d$ and the WGM wavenumber $k_{\mathrm{wg}}$ are varied. The white point denotes the proposed parameters of the experiment, giving $\gamma_{f}=0.035\Gamma_{0}$.
Although a disordered atomic ensemble can become subradiant to free space through a steady state excitation, the degree of superradiant interaction with the cavity will also decrease, resulting in $\gamma_{c}/\gamma_{f}\sim NC_{1}$, similar to the superradiant scaling found in the TDS. The TDS, on the other hand, is more useful in utilizing the full superradiance benefit of this platform. In this section, we explore the potential of atom arrays in reaching subradiance in free space through the simple TDS excitation.
The next stage of the experiment involves the implementation of an array of atoms trapped above the ring resonator. According to Chang et al. (2019), the atoms will be trapped using two-color evanescent fields, resulting in the atoms being separated by about $0.3\lambda$, where $\lambda$ is the resonant wavelength of the light. This is close to half the wavelength of the high-k excitation mode of the ring resonator. When the atoms are excited through the resonator, it can facilitate almost perfectly subradiant states into free space due to the alternating phases in adjacent atoms.
The higher wavenumber of the waveguide excitation $k_{\mathrm{wg}}$ plays a major role in contributing to accessing highly subradiant states in the case of atom arrays. Experimentally, this is only dependent on the effective refractive index $n_{\mathrm{eff}}=k_{\mathrm{wg}}/k_{0}$ of the WGM. The decay rate into free space as a function of $k_{\mathrm{wg}}$ and the separation of the atom $d$ has been depicted in Fig. 5 for a perfect atom array with no disorder. Although very good subradiance ($\gamma_{f}<0.1\Gamma_{0}$) can only be achieved with $n_{\mathrm{eff}}>1$ and separations $d<0.5\lambda$, there are still other regions where nominal subradiance ($\gamma_{f}<\Gamma_{0}$) can be achieved. As mentioned at the end of Sec. II.5, in the low-intensity limit, $\gamma_{f}$ is only dependent on the relative phase between the atomic excitation and is therefore independent of the single atom cooperativity $C_{1}$.
The white dot in Fig. 5 marks the proposed parameters of the upcoming experiment. For a perfect array with $N=20$, separation of $d=0.3\lambda$, and effective refractive index $n_{\mathrm{eff}}=1.69$, the free space decay rate $\gamma_{f}=0.035\Gamma_{0}$.
III.3.1 Effects of Disorder
Although atom arrays show great promise in achieving subradiance in free space, inevitable defects due to experimental implementations can be limiting. We study how two common sources of imperfection can disturb the delicate coherences built in an ideal subradiant state.
The coupling of the cavity mode with the atoms has an exponential dependence on the z-position due to the nature of the evanescent coupling. Thus, the randomness in the z-position due to trapping will cause a stochastic non-uniformity in $C$ of each atom. This randomness in the single atom cooperativity can be treated as a source of decoherence.
In Fig. 6(a), the decay rate into free space is plotted as a function of the r.m.s. size $\delta z$ of the atomic wavepacket along the z-axis. In this case, the spread in the other two directions is assumed to be negligible to isolate the impact of $\delta z$. The red squares denote the case when all the atoms have a uniform $C$, and the blue circles denote the situation where the $C$ is z-position dependent. Just the spread in position alone can cause decoherence and reduce the degree of subradiance, but the stochastic non-uniformity in $C$ causes a much larger effect. At $\delta z=50\mathrm{nm}$ spread, the reduction in subradiance due to the non-uniformity in $C$ is around $8$ times larger than the reduction in the case with uniform $C$. This emphasizes the importance of confining the atoms tightly in the z-direction in the case of a system with evanescent coupling.
In Fig. 6(b), the effect of the filling fraction of the array is depicted. We grow the system size by probabilistically placing one atom in each added site based on the filling fraction until the system contains 20 or 40 atoms. Although the number of array sites will vary with the filling fraction, this is adopted to maintain the scaling of the waveguide interaction with the number of atoms. The trend is almost linear, connecting the points from $\Gamma_{0}$ asymptotically at $0\%$ filling fraction to the minimum possible decay rate at $100\%$ filling fraction. This emphasizes the importance of implementing arrays with a high filling fraction to achieve good subradiance in free space. Even if the state is completely dark for a perfect array, a $50\%$ filling fraction can only allow a minimum of $\gamma_{f}=0.5\Gamma_{0}$.
Figure 6: Effects of disorder in atom arrays with the TDS. The decay rate into free space has been plotted versus (a) the spread $\delta z$ along the z-axis and (b) the filling fraction.
III.3.2 Circular Array
Figure 7: Decay rate of the TDS from an ordered atom array versus the number of atoms in the array. Red squares denote a linear array, and blue circles denote a ring-shaped array.
For the linear array that we have been discussing so far, most of the emission into free space occurs at the ends of the array. Reference Asenjo-Garcia et al. (2017) has shown an exponential improvement in subradiance when using a circular array configuration. The structure of the microring resonator naturally lends itself to the potential of having a ring of atoms trapped along the resonator.
Figure 7 shows the comparison between a linear array and a ring-shaped array trapped along the microring resonator. It depicts the decay rate into free space that is achievable using the TDS excitation through the waveguide as a function of the number of atoms in the array. This is similar to Fig. 7 in Ref. Asenjo-Garcia et al. (2017). We can almost reach a perfect subradiant state with the ring resonator setup and evanescent field traps.
A ring of atoms can form its own WGM, resulting in minimal loss of photons into free space. The interactions between this mode and the WGM of the resonator hold great potential for applications in quantum control and photon storage.
Although there is great promise with a ring-shaped array due to the exponential scaling, the effects of disorder described in Sec. III.3.1 will still limit the effectiveness of the system.
IV Conclusion
We have studied the effects of collective dipole-dipole interactions in a system of atoms trapped in the vicinity of a nanophotonic microring resonator. We modeled and simulated the experimental setup in Ref. Zhou et al. (2025) to understand the dynamics and inform potential future research directions. The atoms can interact with each other in two ways: the free space dipole-dipole interaction and the microring cavity interaction. Different properties and the mismatched wavenumber of the two lead to interesting dynamics.
We studied the spectrum and decay rates of a cloud of atoms trapped near the resonator. We explored how the two different types of excitation, using a short pulse (for the timed-Dicke state) or a long pulse (for the steady state), affect the decay dynamics of the system. We also studied how the ratio of the decay rate into the desired and undesired modes scales with the number of atoms.
We explored the implications of trapping atoms in sub-wavelength arrays along the resonator. We studied the effects of the high wavenumber excitation through the microring and explored potential impacts of disorder on subradiance in experimental implementations. This platform also holds the potential to form a ring of trapped atoms along the microring resonator to achieve a much better scaling on subradiance when compared to linear arrays of atoms.
This versatile light-matter interaction platform holds great promise with its many degrees of control and tunable parameters. Implementing atom arrays will open up the possibility of "selective radiance." Expanding the study to three-level atoms opens possibilities to study applications such as photon storage.
Acknowledgements.
The work of D.S. and F.R. was supported by the National Science Foundation under Award No. 2410890-PHY. The work of X.Z. and C.L.H. was supported by the AFOSR (Grant NO. FA9550-22-1-0031) and the ONR (Grant NO. N000142412184). This research was supported in part through computational resources provided by Information Technology at Purdue University, West Lafayette, Indiana.
Appendix A Adiabatic Elimination of the cavity
The cavity has two sources of loss. It can intrinsically lose a photon with the rate $\kappa_{i}$. Cavity photons can also hop to the bus waveguide with coupling rate $\kappa_{e}$. This cavity decay is defined by the Lindblad operator,
$\mathcal{L}_{R}(\hat{\rho})=(\kappa_{i}+\kappa_{e})[\hat{a}\rho\hat{a}^{\dagger}-\frac{1}{2}\hat{a}^{\dagger}\hat{a}\rho-\frac{1}{2}\rho\hat{a}^{\dagger}\hat{a}].$ (27)
By enforcing that the cavity reaches a steady state at every time-step of the internal dynamic evolution ($\langle\dot{\hat{a}}\rangle=0$), we arrive at an expression for $\langle\hat{a}\rangle$ depending on the atomic excitations.
$\langle\hat{a}\rangle=\frac{1}{\tilde{\kappa}}\bigg{(}\sum_{j}g_{j}e^{i\phi_{j}}\langle\hat{\sigma}_{j}^{-}\rangle+\eta\bigg{)}.$ (28)
By replacing $\hat{a}$ in Eq. (8) and Eq. (27), the cavity can be adiabatically eliminated, and the master equation can be re-written to arrive at Eqs. (10), (11) and (13).
Appendix B Orthogonality of the eigenmodes
The atom-cavity interaction matrix $\mathbf{G}_{c}$ is skew-symmetric because of the running-wave nature of the resonator. This results in purely imaginary eigenvalues, meaning none of the eigenmodes of the atom-cavity interaction has any energy shift. These eigenmodes also follow the familiar orthogonality relation $\vec{U}\cdot\vec{V}=\vec{U}^{\dagger}\vec{V}$.
On the other hand, the interaction matrix of the free space modes $\mathbf{G}_{f}$ is complex-symmetric. This means that $G^{f}_{ij}=G^{f}_{ji}$. This results in the eigenmodes having a different orthogonality relation $\langle\vec{U},\vec{V}\rangle=\vec{U}\vec{V}$.
The combined matrix will be neither Hermitian nor Complex-symmetric, resulting in different left $(\vec{L})$ and right eigenvectors $(\vec{R})$ and a different orthogonality relation.
$\begin{split}\mathbf{M}\vec{R}_{\alpha}=\lambda_{\alpha}\vec{R}_{\alpha}\\ \vec{L}^{\dagger}_{\alpha}\mathbf{M}=\lambda_{\alpha}\vec{L}^{\dagger}_{\alpha},\end{split}$ (29)
where $\lambda_{\alpha}$ denotes the eigenvalue of eigenmode $\alpha$ of the full interaction matrix. The left and right eigenvectors $\vec{L}_{\alpha}$ and $\vec{R}_{\alpha}$ can then be normalized to follow
$\vec{L}^{\dagger}_{\alpha}\vec{R}_{\alpha^{\prime}}=\delta_{\alpha,\alpha^{\prime}}.$ (30)
Using this orthogonality relation, we can decompose the driving laser term in Eq. (18) into components that each drive a particular eigenmode $\alpha$
$\tilde{\Omega}_{\alpha}=\vec{L}^{\dagger}_{\alpha}\vec{\Omega}.$ (31)
Since each eigenmode evolves individually with its characteristic decay rate, the equation of motion of each eigenmode can be solved separately. The contribution from each eigenmode $w_{\alpha}$ can be calculated for the SS in the low intensity,
$w_{\alpha}^{\mathrm{SS}}=-\frac{\tilde{\Omega}_{\alpha}}{2(\lambda_{\alpha}-\Delta_{A})},$ (32)
and for the TDS
$w_{\alpha}^{\mathrm{TDS}}\propto\tilde{\Omega}_{\alpha},$ (33)
with the proportionality depending on the pulse area. These can be used with Eq. (20) to calculate the final state and the composition with respect to eigenmodes.